if the average income rises from 18000 per year to 22000 per year annual gasoline consumption per household rise from 1000 to 1500 gallons the income elasticity of demand for gas is
1
Expert's answer
2020-09-22T15:27:22-0400

"IncomeElasticityOfDemand=\frac{\% in\ quantity\ demanded}{\% in\ income}" "IncomeElasticityOfDemand=\frac{(1500-1000)/1000*100\%}{(22000-18000)/18000*100\%}\frac{50\% }{22.2\%}=2.25"